OpenStudy (anonymous):

Given f(x) = 4x+1 over 3, solve for f^-1(3) A. 2 B. 4 C. 6 D. 8

OpenStudy (anonymous):

Please help !!

OpenStudy (anonymous):

let's call f(x) like y. So we get: y=4x+1 now soolve for x. x=(y-1)/4 this is f^-1(y) for y=3 x=(3-1)/4=?

OpenStudy (anonymous):

ups, didnt notice the over 3 part

OpenStudy (anonymous):

let's call f(x) like y. So we get: y=(4x+1)/3 now soolve for x. x=(3y-1)/4 this is f^-1(y) for y=3 x=(3*3-1)/4=?

OpenStudy (anonymous):

2 so that means the answer is a right ?

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

yes @Niyah98

OpenStudy (anonymous):

Thank you @myko